OpenStudy (poisoned_ruby):

the charge on a capacitor in a circuit containing C,a resistance R,and a source of voltage E is given by q=CE(1-e^(-t/RC)).show that this satisfies the equation Rdq/dt+q/C=E

OpenStudy (michele_laino):

Hint: If I compute the first derivative of the function \(\Large q(t)\), then I get: \[\Large \frac{{dq}}{{dt}} = EC \cdot \frac{1}{{RC}}\exp \left( { - \frac{t}{{RC}}} \right)\] please substitute that first derivative, together the same function \(\Large q(t)\) into your differential equation, you should get an identity

OpenStudy (poisoned_ruby):

Is that exp the same as e?

OpenStudy (poisoned_ruby):

e as in the opposite of a natural log?

OpenStudy (irishboy123):

yes

OpenStudy (poisoned_ruby):

Thank you :)

OpenStudy (irishboy123):

no, thank "the maestro"

OpenStudy (poisoned_ruby):

Thank both of you lol. You helped too :) I had no idea what exp meant. Shows how much I know about claculus :/

OpenStudy (poisoned_ruby):

To find that derivative, did you first distribute the CE and then use d/dx e^x=e^x * x^i?

pooja195 (pooja195):

@IrishBoy123 @Michele_Laino

OpenStudy (irishboy123):

hey ruby, i think we started with: \[\Large q(t)=CE(1-e^{\frac{-t}{RC}})\] and we have to show that \(\Large R\dot q+\dfrac{q}{C}=E\) is true 

OpenStudy (poisoned_ruby):

Yes I'm just not sure how we got to

OpenStudy (poisoned_ruby):

I'm sorry if I seem incompetent lol.

OpenStudy (irishboy123):

ruby, are you online now???

OpenStudy (poisoned_ruby):

Yes

OpenStudy (irishboy123):

cool did we get the first bit right: \(\large q(t)=CE(1-e^{\frac{-t}{RC}})\) ??

OpenStudy (poisoned_ruby):

Yes, and you're correct in stating that we need to prove that \[q=CE(1-e ^{-t/RC}) \] satisfies the second equation

OpenStudy (irishboy123):

so we need : \(\large \dot q(t)=CE \dfrac{d}{dt}\left\{(1-e^{\frac{-t}{RC}})\right\}\)

OpenStudy (poisoned_ruby):

Would you distribute the CE and then use the chain rule theorem?

OpenStudy (irishboy123):

the CE is a constant. keep it outside.

OpenStudy (poisoned_ruby):

Alright, so just use chain rule on e^-t/RC?

OpenStudy (irishboy123):

\(\large \dfrac{d}{dt}(1-e^{\frac{-t}{RC}}) = \dfrac{d}{dt}(1)-\dfrac{d}{dt}e^{\frac{-t}{RC}}) \)

OpenStudy (irishboy123):

\(\large \dfrac{d}{dt}(e^{-\frac{t}{RC}}) = -\frac{1}{RC}~e^{\frac{-t}{RC}}\)

OpenStudy (irishboy123):

ruby keep asking, if the answers are not helpful. :p

OpenStudy (poisoned_ruby):

And then I substitute that into the second equation? And try to get E=E?